OpenStudy (anonymous):

Can someone explain to me the limitations on the base of an exponential function?

OpenStudy (anonymous):

Base should be positive. f(x)=a^x, for negative a the function is not defined. for a= 0 your function is not defined at x<=0.

OpenStudy (anonymous):

that was an example ^

OpenStudy (anonymous):

Should "always" be positive? Or just most of the time.

OpenStudy (anonymous):

?*

OpenStudy (helder_edwin):

Yes \[ \large a^x=e^{x\ln a} \]

OpenStudy (anonymous):

Thanks